How many atoms are there in 5 moles of phosphorus
1
Expert's answer
2021-08-29T16:55:02-0400

Given:

"n=5\:\rm mol"


The amount of moles is given by

"n=\frac{N}{N_A}"

Hence, the number of atoms is

"N=nN_A\\
=5\:\rm mol*6.02*10^{23}\: mol^{-1}=3.01*10^{24}"